OpenStudy (anonymous):

What is the solution to the rational equation 3 over x minus 2 = 2 over x minus 3 ?

OpenStudy (mertsj):

\[\frac{3}{x-2}=\frac{2}{x-3}\]

OpenStudy (mertsj):

Cross multiply and solve.

OpenStudy (anonymous):

x = 0?

OpenStudy (mertsj):

What is 3(x-3)?

OpenStudy (anonymous):

distribute right ??

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

ohh i got it

OpenStudy (anonymous):

3x-9

OpenStudy (mertsj):

What is 2(x-2)?

OpenStudy (anonymous):

2x-4

OpenStudy (mertsj):